### v1

# Approaching the Quantum Speed Limit in Quantum Gates with Geometric Control

Abstract
We present a geometric optimization method for implementing quantum gates by optimally controlling the Hamiltonian parameters, aiming to approach the Mandelstam-Tamm Quantum Speed Limit (MT-QSL). Achieving this bound requires satisfying precise geometric conditions governing the evolution of quantum states. We extend this geometric framework to quantum unitary operators in arbitrary dimensions and analyze the conditions necessary for saturation of the bound. Additionally, we demonstrate that the quantum brachistochrone, when generalized to operators, does not generally reach the MT-QSL bound. Finally, we propose a systematic optimal control strategy based on geometric principles to approach the quantum speed limit for unitary driving.
Approaching optimality in quantum driving is both a challenging and crucial task for quantum information systems. Reducing evolution time while maintaining constant resources enhances efficiency and can potentially increase the resilience of quantum protocols against external perturbations. The concept of the Quantum Speed Limit (QSL), which defines the minimal evolution time required for an initial state to reach an orthogonal subspace, has been a subject of growing interest and extensive investigation [1, 2, 3, 4, 5, 6, 7, 8], with the recent generalization to open  [9, 10, 11, 12], non-Markovian [13] and non-linear [14, 15] quantum systems. Among recent developments, we highlight the search for unified geometric definitions of the QSL [16], the investigation of the relation of the QSL with the system quantumness [17, 18], with interference effects [19, 20], or with statistical properties [21, 22] of the system, and finally an effort to approach the QSL in concrete experimental setups [23]. The QSL is deeply linked to the geometry of quantum state space [24] and involves identifying quantum trajectories that correspond to geodesics in the Fubini-Study metric. When formulated as a quantum brachistochrone problem, it allows for analytical [25, 26] or approximate [27] solutions for specific classes of Hamiltonians. However, optimal solutions to the quantum brachistochrone problem do not always saturate the QSL bounds associated with the evolution of a single quantum state [28]. While a significant body of literature has focused on the QSL for the evolution of individual states, the exploration of QSLs governing the evolution of sets of quantum states remains in its early stages [29, 30, 31, 32].
The goal of this paper is to present an alternative derivation of such bounds, inspired by the approach of L. Vaidman [3], which also introduces a geometric optimization method for implementing time-optimal unitary quantum gates. More specifically, we identify two necessary criteria for optimal driving: one involves a phase condition, while the other has a geometric interpretation. Indeed, the deviation from the optimal Mandelstam-Tamm QSL bound increases as the time evolution of the operator deviates from a specific time-dependent plane in the Hilbert space of quantum operators.
In Sec. I, we generalize the Mandelstam-Tamm bound on the quantum speed limit using an approach similar to that developed by L. Vaidman [3]. We then discuss the generalization of the Margolus-Levitin bound for time evolution operators. In Sec. II, we explore the conditions under which the bound is saturated, illustrating these with various examples, and analyze the quantum brachistochrone from the QSL perspective. Finally, in Sec. III, we use the geometric framework to identify the additional constraints that must be incorporated into an optimal control approach to achieve the maximal Quantum Speed Limit (QSL) for a given unitary operator. As intuitively expected, the efficiency of optimized quantum driving relative to the QSL is shown to depend on the degree of control over the Hamiltonian parameters. To illustrate this, we consider two scenarios using a concrete example of unitary gates in a four-dimensional space: one with full control over the Hamiltonian parameters, and another with more restricted control.
I Quantum Speed Limits for unitary operators
In this paragraph, we derive the two standard bounds for the quantum speed limit in the context of quantum control of time evolution operators. We consider a $d$-dimensional Hilbert space $\mathcal{H}$ and the associated space of linear operators $\mathcal{L}(\mathcal{H})$, on which we define the scalar product $\langle U_{1},U_{2}\rangle=\frac{1}{d}{\rm Tr}\left[U_{1}^{\dagger}U_{2}\right]$. Unitary operators $U$ have unit norm $\langle U,U\rangle=1$.
I.1 Mandelstamm and Tamm bound for quantum operators
Using geometric arguments, L. Vaidman derived a minimal time required to drive a given initial quantum state to any orthogonal quantum state  [3]. Following a similar approach, we derive a QSL for the full quantum evolution operator and establish a lower bound on the evolution time required to reach an orthogonal quantum operator.
Given an arbitrary, possibly time-dependent, Hamiltonian $H(t)$ and a unitary evolution operator $U(t)$, we decompose $H(t)U(t)$ along the direction spanned by $U(t)$ and its orthogonal subspace as follows:
$H(t)U(t)=\lambda_{/\!/}(t)U(t)+\lambda_{\perp}(t)U_{\perp}(t)$ (1)
where $U_{\perp}(t)$ is a linear operator such that $\langle U(t),U_{\perp}(t)\rangle=0$ and $\langle U_{\perp}(t),U_{\perp}(t)\rangle=1$ (note that $U_{\perp}(t)$ is in general non-unitary). By projecting both sides of the equation onto the operator $U(t)$, we obtain $\lambda_{/\!/}(t)=\frac{1}{d}{\rm Tr}[U^{\dagger}(t)H(t)U(t)]=\frac{1}{d}{\rm Tr}[H(t)]=\langle H(t)\rangle$ where we define $\langle O\rangle=\frac{1}{d}{\rm Tr}[O]$ for any operator $O$. Next, we take the scalar product with the Hermitian conjugate $H(t)U(t)$ on both sides of Eq. (5), and readily obtain $\frac{1}{d}{\rm Tr}[H(t)^{2}]=|\lambda_{/\!/}(t)|^{2}+|\lambda_{\perp}(t)|^{2}$ where we have used $H(t)^{\dagger}=H(t)$, $\langle U(t),U(t)\rangle=\langle U_{\perp}(t),U_{\perp}(t)\rangle=1$ and $\langle U_{\perp}(t),U(t)\rangle=\langle U(t),U_{\perp}(t)\rangle=0.$ As a result, we find $|\lambda_{\perp}(t)|^{2}=\langle H(t)^{2}\rangle-\langle H(t)\rangle^{2}$, i.e. $|\lambda_{\perp}(t)|=\Delta H(t).$ Consequently, the operator $U_{\perp}$ can be expressed as
$U_{\perp}(t)=\frac{1}{\Delta H(t)}\left(H(t).U(t)-\langle H(t)\rangle U(t)\right)$ (2)
as long as $\Delta H(t)\neq 0$.
Using the decomposition above, the Schrödinger equation for the evolution of $U(t)$ reads:
$\dot{U}=-\frac{i}{\hbar}\left(\langle H\rangle\>U+\Delta H\>U_{\perp}\right).$ (3)
The initial evolution operator is the identity matrix $U_{0}=\mathbb{I}$. The time derivative of its overlap with the evolved quantum operator can be estimated as:
$\frac{d|\langle U_{0},U(t)\rangle|^{2}}{dt}=\frac{-2\Delta H(t)}{\hbar}{\rm Re}\left[(-i)\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle\right]$
We have omitted the term proportional to the average energy, which corresponds to a purely imaginary number. The initial operator $U_{0}$ can be decomposed as
$U_{0}=\langle U(t),U_{0}\rangle U(t)+\langle U_{\perp}(t),U_{0}\rangle U_{\perp}(t)+U_{\perp\perp}(t)$ (5)
where $U_{\perp\perp}(t)$ is a linear operator that is orthogonal to both $U(t)$ and $U_{\perp}(t)$. We obtain
$|\langle U_{0},U_{\perp}(t)\rangle|$ $=$ $\sqrt{1-|\langle U_{0},U(t)\rangle|^{2}-||U_{\perp\perp}(t)||^{2}}$ (6)
$\leq$ $\sqrt{1-|\langle U_{0},U(t)\rangle|^{2}}.$
where the inequality is saturated if and only if
$||U_{\perp\perp}(t)||^{2}=0$ (7)
i.e. if $U_{0}$ belongs to the plane spanned by $\{U(t),H(t)U(t)\}$ at all times.
Combining Eq. (I.1) with the inequality (6), we obtain:
$\frac{d|\langle U_{0},U(t)\rangle|^{2}}{dt}\leq\frac{2\Delta H(t)}{\hbar}|\langle U_{0},U(t)\rangle|\sqrt{1-|\langle U_{0},U(t)\rangle|^{2}}$ (8)
By setting $\cos\phi=|\langle U_{0},U(t)\rangle|$, the inequality (8) can expressed as an upper bound on the time derivative of the angle $\phi$:
$|\dot{\phi}|\leq\frac{\Delta H}{\hbar}$ (9)
To saturate this QSL, two conditions must be simultaneously fulfilled
$|\langle U_{0},U(t)\rangle|^{2}+|\langle U_{0},U_{\perp}(t)\rangle|^{2}$ $=$ 11 11
${\rm Re}\left[\langle U(t),U_{0}\rangle\langle U_{0}(t),U_{\perp}(t)\rangle\right]$ $=$ 00 0 (10)
The first condition states that $U_{\perp\perp}(t)=0,$ meaning that the time derivative $\dot{U}(t)$ must always lie in the plane spanned by $\{U_{0},U(t)\}$. Since the evolution operator $U(t)$ is also required to remain unitary at all times, this imposes a significant constraint on the possible quantum paths that saturate the QSL bound. In practice, it is not guaranteed that such a path exists in Hilbert spaces with dimensions larger than two. In such cases, one can quantify the deviation from the QSL when these conditions are not satisfied. Setting $\alpha(t)=||U_{\perp\perp}(t)||^{2}=\sin^{2}\phi(t)-|\langle U_{0},U_{\perp}(t)\rangle|^{2}$ and $\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle=|\langle U(t),U_{0}\rangle\langle U_{0},U_{\perp}(t)\rangle|e^{i\beta(t)}$, one has
$\dot{\phi}=\frac{\Delta H(t)}{\hbar}\sqrt{1-\frac{|\alpha(t)|^{2}}{\sin\phi(t)^{2}}}\sin\beta(t).$ (11)
The ideal case of QSL saturation corresponds to $\alpha(t)=0$ (plane condition) and to $\beta(t)=\pm\frac{\pi}{2}$ (phase condition) at all times. An optimization method based on these two relations will be presented in Sec. III.
I.2 Margolus-Levitin quantum speed limit
Interestingly, one can also retrieve the Margolus-Levitin bound [33] in this operatorial formalism. For this purpose, we consider a time-independent Hamiltonian $H_{0}=\sum_{n=1}^{d}E_{n}|n\rangle\langle n|$ with $E_{1}\leq E_{2}\leq...\leq E_{d}$ with eventually degenerate eigenvalues. The associated evolution operator reads $U(t)=\sum_{n=1}^{d}e^{-\frac{i}{\hbar}E_{n}t}|n\rangle\langle n|,$ and one finds
$D(t)=\frac{1}{d}{\rm Tr}\left[U_{0}^{\dagger}U(t)\right]=\frac{1}{d}\sum_{n=1}^{d}e^{-\frac{i}{\hbar}E_{n}t}$ (12)
where we have considered the initial evolution operator $U_{0}=\mathbb{I}$ (the same argument could be applied to a different gate $U_{f}$). One can proceed in the same way as in the proof for single states (with here $|c_{n}|^{2}\equiv 1/d$). Thus, we have that
${\rm Re}\left[D(t)\right]$ $=$ $\frac{1}{d}\sum_{n=1}^{d}\cos\left(\frac{E_{n}}{\hbar}t\right)$ (13)
$\geq$ $\frac{1}{d}\sum_{n}\left[1-\frac{2}{\pi}\left(\frac{E_{n}t}{\hbar}+\sin\left(\frac{E_{n}t}{\hbar}\right)\right)\right]$
$\geq$ $1-\frac{2}{\pi}\langle H_{0}\rangle\frac{t}{\hbar}+\frac{2}{\pi}{\rm Im}[D(t)]$
where we have used that
$\cos(x)\geq 1-\frac{2}{\pi}\left(x+\sin(x)\right)$ (14)
valid for $x\geq 0$, which becomes an equality if and only if $x=0,\pi$. Assuming that $D(\tau)=0$, i.e., that the evolution leads to an orthogonal gate after time $\tau$, the previous inequality gives a lower bound on this duration
$\tau\geq\frac{\pi\hbar}{2\langle H_{0}\rangle}$ (15)
This QSL bound is valid for any constant and positive Hamiltonian $H_{0}$. The replacement $H_{0}^{\prime}=H_{0}-E^{\prime}$ merely adds a global phase factor and thus leaves invariant the time $\tau$. As long as $E^{\prime}\leq E_{1}$ to guarantee the positivity of the Hamiltonian, the argument above yields a similar bound with $\langle H_{0}^{\prime}\rangle=\langle H_{0}\rangle-E^{\prime}$ instead of $\langle H_{0}\rangle.$ The tighter bound is obtained with $E^{\prime}=E_{1},$ so that one can replace $\langle H_{0}\rangle$ by $\langle H_{0}\rangle-E_{1}$ in the denominator of Eq. (15).
One might wonder under which conditions the QSL bound is attained. To simplify the discussion, we consider the Hamiltonian $H_{0}^{\prime}=H_{0}-E_{1}$ with a ground state of zero energy, and the associated quantity $D(t)$. All the terms in the sum of Eq. (13) must saturate the inequality Eq. (14). The first term, by construction, corresponds to $x=0$, and there can be at most one additional term corresponding to $x=(E_{j}-E_{1})\tau/\hbar=\pi$ for some $j>1$. The only way to saturate the bound is therefore to choose the evolution time as $\tau=\pi/\hbar(E_{j}-E_{1}),$ and to have a Hamiltonian of the form:
$H_{0}=\sum_{n\in\mathcal{S}_{1}}E_{1}|n\rangle\langle n|+\sum_{n\in\mathcal{S}_{j}}E_{j}|n\rangle\langle n|$ (16)
where $\{\mathcal{S}_{1},\mathcal{S}_{j}\}$ is a partition of the eigenstate basis in two degenerate subspaces of energies $\{E_{1},E_{j}\}$.
The coincidence of the MT and ML QSL bounds occurrs when $\Delta H=\langle H\rangle-E_{1}$. Since $\langle H\rangle-E_{1}=\frac{n}{d}(E_{j}-E_{1})$ and $\Delta H=\frac{1}{d}\sqrt{n(d-n)}(E_{j}-E_{1})$ (with $n={\rm card}\mathcal{S}_{j}$), this is simply equivalent to $n=d/2$. The two bounds coincide and saturate when the quantum system exhibits dynamics essentially identical to that of a two-level system.
II Saturation of the MT-QSL bound for specific families of unitary dynamics
To illustrate the constraints imposed by the Mandelstam-Tamm QSL for unitary operators, we examine the potential saturation of this quantum speed limit in low-dimensional systems and discuss the brachistochrone approach within the QSL framework.
II.1 Saturation of the MT-QSL in a 2D system
In a two-dimensional system, up to an arbitrary phase, a generic quantum state is fully defined by two angles, in agreement with the conventional Bloch sphere representation. As a result, when defining a general unitary operator, the mapping of the first state of the canonical basis corresponds to two degrees of freedom. The second basis state is necessarily mapped to an orthogonal state, but its phase can still be chosen independently. Therefore, quantum gates in a two-dimensional Hilbert space are parametrized by three degrees of freedom.
First, we examine the saturation of the QSL within a family of unitary dynamics parameterized by two control parameters $U(\theta(t),\varphi(t))=e^{-\frac{i}{2}\varphi(t)}|\psi_{+}(\theta(t))\rangle\langle 0|+e^{\frac{i}{2}\varphi(t)}|\psi_{-}(\theta(t))\rangle\langle 1|$ with the two orthogonal states $|\psi_{+}(\theta)\rangle=\cos(\theta/2)|0\rangle+\sin(\theta/2)|1\rangle$ and $|\psi_{-}(\theta)\rangle=-\sin(\theta/2)|0\rangle+\cos(\theta/2)|1\rangle$. This corresponds to a subset of the possible unitary gates in a two-dimensional system, but the extension to the more general case, which includes all possible unitary operators, is straightforward. Here, we aim to determine the dynamics of the parameters $(\theta(t),\varphi(t))$ such that the QSL is saturated at all time. The Hamiltonian is determined by $H(t)=i\hbar U^{\dagger}(t)\dot{U}(t)$ and is given by $H(t)=\frac{\hbar}{2}\left(\dot{\theta}\sigma_{y}+\dot{\varphi}\sigma_{z}\right)$ where $\sigma_{x,y,z}$ are the usual Pauli operators. We obtain $\Delta H=\frac{\hbar}{2}\sqrt{\dot{\theta^{2}}+\dot{\varphi}^{2}}.$ On the other hand, the overlap with the initial evolution operator is given by $\langle U_{0},U(t)\rangle=\cos\left(\frac{1}{2}\theta(t)\right)\cos\left(\frac{1}{2}\varphi(t)\right)$. By setting $\cos\phi=\langle U_{0},U(t)\rangle$, we find
$\dot{\phi}=\frac{\dot{\theta}\sqrt{1-x^{2}}y+\dot{\varphi}x\sqrt{1-y^{2}}}{2\sqrt{1-x^{2}y^{2}}}$ (17)
with $x=\cos\left(\frac{\theta(t)}{2}\right)$ and $y(t)=\cos\left(\frac{\varphi(t)}{2}\right).$ As $|y|\leq 1$, one finds $\sqrt{1-x^{2}}y=\sqrt{y^{2}-x^{2}y}\leq\sqrt{1-x^{2}y^{2}}$ and similarly $x\sqrt{1-y^{2}}\leq\sqrt{1-x^{2}y^{2}}$, which yields:
$\dot{\phi}\leq\frac{\dot{\theta}+\dot{\varphi}}{2}$ (18)
One has $\dot{\theta}+\dot{\varphi}\leq\sqrt{\dot{\theta}^{2}+\dot{\varphi}^{2}}$ by concavity of the function $x\rightarrow\sqrt{x^{2}+y^{2}}$ for any $y$, so that our result is consistent with the MT-QSL (9). Saturation occurs only if $x=0$ or $y=0$ at all times, meaning that either $\dot{\theta}=0$ at all times, or $\dot{\varphi}=0$ at all times. The former case corresponds to a trivial scenario where the unitary operator remains the identity throughout. Therefore, we must have $\dot{\varphi}=0$ at all times. We recover here the fact that the geodesics on a sphere coincide with its great arcs. Note that one may alternatively define an optimal QSL with respect to some unitary operator $U_{1}$ reached after a certain evolution. This example shows that saturating the QSL severely constrains the choice of possible quantum paths. To achieve QSL saturation, only one of the two angles $(\theta(t),\varphi(t))$ can evolve at a time.
II.2 Investigation of the QSL limit in 3-level systems with Rodrigues formula
In Ref. [34], a method was proposed to control the dynamics of a three-level system by separating the evolution into two components: population changes, parameterized using Rodrigues’ rotation formula, and phase changes. This approach was used to inversely design the Hamiltonian of the three-level system, enabling the desired transition while ensuring allowed couplings and suppressing forbidden ones.
Similarly, we focus here on a three-state system and the family of unitary operators ${U_{r}(t)}$ that govern population transfer. Specifically, we consider mappings between quantum states with positive real-valued coefficients, noting $\mathcal{S}=\{|\psi\rangle|\forall n\in\{1,2,3\}\>\langle n|\psi\rangle\in\mathbf{R}^{+}\}$ the corresponding set of quantum states. In other words, if $|\psi(0)\rangle=a_{1}(0)|1\rangle+a_{2}(0)|2\rangle+a_{3}(0)|3\rangle$, where $a_{n}(0)\in\mathbb{R}^{+}$ for $n\in\{1,2,3\}$, then at any time $|\psi(t)\rangle=U_{r}(t)|\psi(0)\rangle$, the condition $\langle n|\psi(t)\rangle\in\mathbb{R}^{+}$ holds. Owing to their positivity, the unitary operators $U_{r}(t)$ correspond to rotations in a three-dimensional Euclidean space [34]. A 3D vector rotation is described by the following transformation:
$\mathbf{v}^{\prime}=\mathbf{v}\cos\theta+\mathbf{u}(\mathbf{u}\cdot\mathbf{v})(1-\cos\theta)+(\mathbf{u}\times\mathbf{v})\sin\theta\,$ (19)
where $\mathbf{v}$ is an arbitrary vector to which the transformation is applied, $\mathbf{v}^{\prime}$ its transformed image, $\mathbf{u}$ the rotation axis and $\theta$ the angle of rotation. The corresponding relation for the quantum evolution operator is given by
$U(t)=U_{r}(t)=\cos\theta\mathbb{I}+(1-\cos\theta)|u\rangle\langle u|+V\sin\theta$ (20)
where we have introduced $|u\rangle=\sum_{m}u_{m}|m\rangle$ $u$ is a real-valued vector of positive coefficient and unit norm - and $V=\sum_{i,j,k}\epsilon_{ijk}u_{i}|k\rangle\langle j|$, with $\epsilon_{ijk}$ the Levi-Civita symbol. Any evolution operator $U_{r}(t)$ that leaves invariant the subset $\mathcal{S}$, i.e. such that $U_{r}(t)(\mathcal{S})\subset\mathcal{S}$, can be expressed in this form.
Using this formalism, we investigate the saturation of the QSL. One of the conditions (7) indicates that the three operators ${U_{0},U(t),H(t)U(t)=i\dot{U}(t)}$ must lie within a common plane. Expressing $H(t)U(t)$ as $H(t)U(t)=\lambda_{1}(t)\mathbb{I}+\lambda_{2}(t)U(t)$ with $(\lambda_{1}(t),\lambda_{2}(t))\in\mathbb{C}^{2}$, we derive an alternative expression for the Hamiltonian: $H(t)=\lambda_{1}(t)\mathbb{I}+\lambda_{2}(t)U^{\dagger}(t)$. The Hermiticity condition of the Hamiltonian then requires that $U^{\dagger}(t)+U(t)=\lambda(t)\mathbb{I}$ for some complex number $\lambda(t)$. Since $V^{\dagger}=-V$, from Eq. (20) this condition simplifies to $\cos\theta\mathbb{I}+(1-\cos\theta)|u\rangle\langle u|=\lambda_{0}(t)\mathbb{I}$. Clearly, this is only satisfied if $\theta=0$, implying that no motion occurs.
Therefore, it is impossible to saturate the QSL in a three-level system when the evolution operators are restricted to the family ${U_{r}(t)}$, corresponding to unitary matrices with positive real-valued coefficients. Here, we illustrate that the bound may be unattainable. In this context, it is valuable to have a method for quantifying and optimizing the distance to speed optimality, as detailed in Sec. III.
II.3 Quantum Brachistochrone for unitary evolution and the MT-QSL limit
For state-to-state transformations, an evolution under the quantum brachistochrone does not necessarily saturate the QSL [28]. Hereafter, we investigate the relationship between time-optimal unitary evolution [26] and the framework of the operatorial quantum brachistochrone [25].
In this framework, the Fubini-Study metric for quantum operators is utilized. It is defined as $ds_{U}=\langle dU,(1-P_{U})dU\rangle$, where the projection operator $P_{A}(X)=\langle A,X\rangle A$ is introduced for any operators $A$ and $X$. For time-independent Hamiltonians, the condition for the geodesic simplifies to: [26]
$\frac{d}{dt}\left[(1-P_{1})\left(\frac{dU}{dt}U^{\dagger}\right)\right]=0$ (21)
with $P_{1}(X)=\langle\mathbb{I},X\rangle\mathbb{I}=\frac{1}{d}{\rm Tr}[X]\mathbb{I}.$ This condition can be interpreted as the existence of a time-independent operator $M_{0}$ and a time-dependent coefficient $\lambda(t)$ such that
$\dot{U}=M_{0}U(t)+\lambda(t)U(t).$ (22)
As a matter of fact, this condition simply corresponds to a time-independent Schrödinger equation and is less stringent than the plane condition that arises for the MT-QSL bound, which imposes that $\dot{U}=\lambda_{1}(t)\mathbb{I}+\lambda_{2}(t)U(t)$.
Indeed, it is worth noticing that several examples of time-optimal unitary driving in a four-state systems, derived within the brachistochrone framework, are reported in Ref. [26] do not saturate the MT-QSL bound.
In the following example, we consider two interacting spins $1/2$, such that the system Hamiltonian takes the form $H(t)=-\sum_{j=x,y,z}J_{j}(t)\sigma_{j}^{1}\sigma_{j}^{2}+\sum_{a=1,2}B_{a}(t)\sigma_{z}^{a}$ with $\sigma^{1}_{j}:=\sigma_{1}\otimes\mathbf{1}_{2},$ $\sigma^{2}_{j}:=\mathbf{1}_{2}\otimes\sigma_{j}$ where $\sigma_{j}$ are the usual Pauli matrices for $j=x,y,z$, and the parameters $J_{j}(t),B_{a}(t)$ are free parameters determined by the brachistochrone equation. The time-optimal realization of a SWAP gate
$U_{\rm SWAP}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{array}\right)$ (23)
[…]
$U_{\rm``QFT^{\prime\prime}}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&i\end{array}\right)$ (24)
we obtain $|\langle\mathbb{I},U_{\rm QFT}\rangle|=1/\sqrt{8}$ yielding a MT-QSL bound $T_{QSL}=\Delta\phi/\omega\simeq 0.385\pi/\omega$. The time corresponding to the brachistochrone is slightly larger, $T_{BRA}=\sqrt{11}\pi/(8\omega)\simeq 0.415\pi/\omega$. These results demonstrate that determining the path with maximum QSL efficiency is a challenging task, as the departure from the QSL typically depends on the constraints imposed on the Hamiltonian as well as on the chosen target quantum gate.
III Quantum optimal driving near the MT-quantum speed limit
To approach the quantum speed limit for unitary operators, we employ the Pontryagin Maximum Principle (PMP) [35] utilizing a cost functional that incorporates the conditions outlined in Eq. (I.1). The Pontryagin principle offers a practical optimization framework to minimize the distance of the time derivative $\dot{U}(t)$ from the plane spanned by the initial evolution operator $U_{0}=\mathbb{I}$ and the current quantum evolution operator $U(t)$.
III.1 Vaidman conditions and definition of the cost functions
We define the two following positive-valued functions:
$f_{\rm Plane}(t)$ $=$ $1-|\langle 1,U(t)\rangle|^{2}-|\langle 1,U_{\perp}(t)\rangle|^{2},$
$f_{\rm Phase}(t)$ $=$ $\Delta H(t)^{2}{\rm Re}\left[\langle U(t),U_{0}\rangle\langle U_{0}(t),U_{\perp}(t)\rangle\right]^{2},$ (25)
that quantifies the proximity of the evolution to the MT-QSL bound. The bound saturation corresponds to $f_{\rm Plane}(t)=0$ and $f_{\rm Phase}(t)=0$ at all times. The condition $f_{\rm Phase}(t)=0$ implies a phase $\beta(t)=\pm\frac{\pi}{2}$, as defined in Eq. (11). While both conditions can be simultaneously and exactly satisfied in two-dimensional quantum systems, achieving an exact solution in higher-dimensional systems depends on the available control parameters in the Hamiltonian. In such cases, a trade-off might be necessary to approximate the QSL bound by minimizing these associated quantities. Identifying the corresponding quantum path is the primary objective of the optimization algorithm described below.
The goal of our quantum protocols is to achieve a target evolution operator under conditions that maximize the proximity of the Quantum Speed Limit (QSL) to its theoretical bound. To this end, we employ the optimal control formalism adapted to quantum systems [36]. It is important to note that this approach guarantees convergence to a local minimum but does not ensure that this minimum is the global one. In the following, we quantitatively assess the distance to the desired final quantum evolution operator by introducing the terminal cost function $\Phi(U(t_{f}),t_{f})=-|\langle U_{\rm target},U(t_{f})\rangle|^{2}$ which decreases when the final evolution operator gets closer to the target. Finally, to incorporate the various constraints, we define the global cost functional as follows:
$J[U(t),H(t)]$ $=$ $\Phi(U(t_{f}),t_{f})$ (26)
[…]
where the integral over $f_{0}(U(t),H(t),t)=p_{1}f_{\rm Plane}(U(t),H(t),t)+p_{2}f_{\rm Phase}(U(t),H(t),t)+p_{3}[t_{f}\Delta H(t)/\hbar]^{2}$ is the running cost. It includes two continuous components linked to the QSL conditions. The relative weights $p_{1},p_{2}$ represent a trade-off between achieving QSL saturation and obtaining a quantum operator close to the target. For $(p_{1},p_{2})=(0,0),$ the QSL optimization is completely desactivated, and the protocol is expected to deliver a final quantum gate with maximal target fidelity. We have added a small penalty term proportional to the Hamiltonian variance to suppress artificially large couplings (in all our numerical examples, we use a very small weight, $p_{3}=10^{-4}$). Identifying the set of quantum gates connected to the identity along a quantum path that closely follows the QSL (e.g., gates accessible via a protocol maintaining an average quantum speed above a certain percentage of the QSL) presents a significant challenge. Our optimization method offers a numerical approach to address this complex problem.
We next reformulate the cost function in terms of real-valued quantities, a necessary step for applying the PMP principle. To achieve this, we decompose the evolution operators and the Hamiltonian into their real and imaginary components, denoted as: $U(t)=U_{R}(t)+iU_{I}(t)$ and $H(t)=H_{R}(t)+iH_{I}(t)$. Using Eq. (2) to express the operator $U_{\perp}(t)$, we find after some basic algebra:
$f_{\rm Plane}(U(t),H(t),t)$ $=$ $1-\frac{1}{d^{2}\Delta H^{2}}\left[\left({\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{R}]\right)^{2}+\left({\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{I}]\right)^{2}\right],$
$-\frac{1}{d^{2}}\left({\rm Tr}[U_{R}]^{2}+{\rm Tr}[U_{I}]^{2}\right)$
$f_{\rm Phase}(U(t),H(t),t)$ $=$ $\frac{1}{d^{4}}\left[{\rm Tr}[U_{I}]{\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]+{\rm Tr}[U_{R}]{\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]\left({\rm Tr}[U_{R}]^{2}+{\rm Tr}[U_{I}]^{2}\right)\right]^{2}.$
III.2 Equations of motion in the Pontryagin formalism
To apply the Pontryagin formalism, we introduce a real-valued matrix, $\mathbf{X}(t)$, associated with the dynamical variable—in this case, the unitary evolution matrix $U(t)$. From now on, we denote:
$\mathbf{X}(t)=\left(\begin{array}[]{c}U_{R}(t)\\ U_{I}(t)\\ \end{array}\right)$ (28)
$\mathbf{X}(t)$ is a $2N\times N$ real-valued matrix. We denote by $\mathbf{x}(t)$ the corresponding $2N^{2}$-dimensional column vector, obtained by arranging the matrix elements into a single column. The elements are ordered alphabetically, with the real parts listed first. For instance, in a 3-dimensional Hilbert space, $\mathbf{x}(t)=(U_{R11}(t),U_{R12}(t),...,U_{R33}(t),U_{I11}(t),U_{I12}(t),...,U_{I33}(t))^{T}$.
Due to the Hermiticity of the Hamiltonian $H(t)$, $H_{R}(t)$ is defined by at most $N(N+1)/2$ independent real parameters, while $H_{I}(t)$ is determined by at most $N(N-1)/2$ independent parameters. The vector $\mathbf{v}(t)$ serves as a control parameter that defines $H(t)$ as a function $H(\mathbf{v}(t))$. In scenarios where all elements of the Hamiltonian can be independently controlled, $\mathbf{v}(t)$ is $N^{2}$-dimensional. However, in most experimental setups where only a subset of Hamiltonian couplings can be independently tuned, the dimension of $\mathbf{v}(t)$ is smaller.
The Schrödinger equation, $i\hbar\dot{U}(t)=H(\mathbf{v}(t))U(t)$, can be reformulated in terms of real-valued matrices as:
$\dot{\mathbf{X}}(t)=\left(\begin{array}[]{c}\dot{U}_{R}\\ \dot{U}_{I}\end{array}\right)=\frac{1}{\hbar}\left(\begin{array}[]{cc}H_{I}(\mathbf{v}(t))&H_{R}(\mathbf{v}(t))\\ -H_{R}(\mathbf{v}(t))&H_{I}(\mathbf{v}(t))\end{array}\right)\mathbf{X}(t)$ (29)
or, more compactly, as a single real-valued vectorial differential equation
$\dot{\mathbf{x}}(t)=f(\mathbf{x}(t),\mathbf{v}(t),t),$ (30)
governed by a vectorial control $\mathbf{v}(t)$. This provides a suitable framework for applying Pontryagin’s Maximum Principle. Note that in this context, $f(\mathbf{x}(t),\mathbf{v}(t),t)$ is a linear function of $\mathbf{x}(t)$, which can be expressed as $f(\mathbf{x}(t),\mathbf{v}(t),t)=M[\mathbf{v}(t)]\mathbf{x}(t)$, where $M[\mathbf{v}(t)]$ is a matrix determined by the Hamiltonian $H(t)$.
The Pontryagin Hamiltonian is given by
$H_{P}(\mathbf{x}(t),\mathbf{p}(t),\mathbf{v}(t),t)$ $=$ $-\frac{1}{t_{f}}f_{0}(\mathbf{x}(t),\mathbf{v}(t),t)$ (31)
$+$ $\mathbf{p}(t)\cdot f(\mathbf{x}(t),\mathbf{v}(t),t),$
where we have introduced the adjoint vector $\mathbf{p}(t)$. The vector $\mathbf{x}(t)$ and its adjoint $\mathbf{p}(t)$ fulfill by construction the Hamilton equations:
$\dot{\mathbf{p}}=-\frac{\partial H_{P}}{\partial\mathbf{x}},\qquad\qquad\dot{\mathbf{x}}=\frac{\partial H_{P}}{\partial\mathbf{p}}.$ (32)
and the boundary condition (for fixed final time $t_{f}$)
$\mathbf{p}(t_{f})=\left.-\frac{\partial\Phi(\mathbf{x}(t),t)}{\partial\mathbf{x}}\right|_{t=t_{f}}.$ (33)
The PMP states that the optimal control parameter $\mathbf{v}^{*}(t)$ is obtained by maximizing the Pontryagin Hamiltonian at any time $t$ along the trajectory $(\mathbf{x}(t),\mathbf{p}(t))$:
$\left.\frac{\partial H_{P}(\mathbf{x}(t),\mathbf{p}(t),\mathbf{v}(t),t)}{\partial\mathbf{v}}\right|_{\mathbf{v}(t)=\mathbf{v}^{*}(t)}=0.$ (34)
Let us now express the dynamical equation for the adjoint $\mathbf{p}(t)$ in a more explicit form. Following the same correspondence between the evolution operator $U(t)$ and its associated column vector $\mathbf{x}(t)$, the column vector $\mathbf{p}(t)$ can be associated with a complex-valued adjoint operator $U^{P}(t)$. It is important to emphasize that this adjoint operator has no physical significance and is not necessarily unitary—it is merely a mathematical object arising from the optimization method. From Eqs. (32), this operator satisfies the dynamical equation (See Appendix):
$\dot{U}^{P}(t)$ $=$ $\frac{1}{t_{f}}\left(\frac{\partial f_{0}(U(t),H(\mathbf{v}(t)))}{\partial U_{R}}+i\>\frac{\partial f_{0}(U(t),H(\mathbf{v}(t)))}{\partial U_{I}}\right)$ (35)
-- $\frac{i}{\hbar}H(\mathbf{v}(t))U^{P}(t).$
When compared to the dynamics of the unitary operator $U(t)$, the adjoint operator includes additional driving terms associated with the derivatives of the cost functions. Finally, we can express the Pontryagin Hamiltonian and the boundary conditions for the adjoint as:
$H_{P}(\mathbf{x}(t),\mathbf{p}(t),t)$ $=$ $-\frac{1}{t_{f}}f_{0}(U(t),H(\mathbf{v}(t)),t),$ (36)
$+\frac{1}{\hbar}\langle U^{P}(t),H(\mathbf{v}(t))U(t)\rangle$
with
$U^{P}(t_{f})=2\langle U_{\rm target},U(t)\rangle U_{\rm target}.$ (37)
With the dynamics of the quantum operator and its adjoint at hand, one can apply a quantum optimal control procedure similar to that described in Refs. [37, 38]. The time evolution over the interval $[t,t_{f}]$ is discretized into $N_{t}$ time bins, each centered at $t_{m}$ for $1\leq m\leq N_{t}$ (in our numerical examples, $N_{t}=200$). Assuming $M$ independent driving parameters in the Hamiltonian, the time-dependent control parameter $[\mathbf{v}(t)]$ is represented as a $N_{t}\times M$-dimensional vector, $[\mathbf{v}(t)]=[v_{1}(t),...,v_{M}(t)]$.
Starting with an initial guess for the control parameter $[\mathbf{v}^{(0)}(t)]$—typically chosen as constant—we proceed iteratively as follows. For a given (generally time-dependent) control parameter $[\mathbf{v}^{(n)}(t)]$, we numerically propagate the evolution operator $U^{(n)}(t)$ over the interval $[t,t_{f}]$. The boundary condition (37) provides the final value of the adjoint operator $U^{P(n)}(t_{f})$.
Next, using Eq. (35), we perform a numerical back-propagation to determine the adjoint operator on the interval $[0,t_{f}]$. At each discrete time $t_{m}$, the Pontryagin Hamiltonian (36) and its gradient with respect to the control parameter are computed. Finally, we update the control parameter as:
$\mathbf{v}^{(n+1)}(t_{m})=\mathbf{v}^{(n)}(t_{m})$ (38)
$+\epsilon\nabla_{\mathbf{v}}H_{P}(U^{(n)}(t_{m}),U^{P(n)}(t_{m}),\mathbf{v},t_{m}))|_{\mathbf{v}=\mathbf{v}^{(n)}(t_{m})}$
where the parameter $\epsilon>0$ is chosen to be sufficiently small. The procedure is then iterated until a satisfactory solution is reached. With an appropriate choice of $\epsilon$, the cost function should decrease over successive iterations. Alternatively, one can use adaptive values $\epsilon\equiv\epsilon^{(n)}$, which are reduced whenever the cost function (26) fails to decrease during a given iteration.
III.3 Optimization in a 4-dimensional space
In this section, we apply our optimization method in a 4-dimensional space. The target is the quantum QFT gate $U_{\rm``QFT^{\prime\prime}}$ of Eq. (24) which is analogous to a QFT gate on two qubits up to a Hadamard transform (see [39]). We consider an evolution governed by the Hamiltonian $H(t)=\sum_{j=1}^{4}\sum_{i<j}\left[\hbar\left(\Omega^{R}_{ij}(t)+i\>\Omega^{I}_{ij}(t)\right)|j\rangle\langle i|+h.c.\right]+\sum_{i=1}^{4}\hbar\Delta_{ii}(t)|i\rangle\langle i|$ over a total time $t_{f}$. We assume full control over all parameters, i.e. 16-dimensional control vector $\mathbf{v}(t)=(\Delta_{ii}(t),\Omega_{ij}^{R}(t),\Omega_{ij}^{I}(t))$, ordered alphabetically for $(i,j)\in\{1,2,3,4\}^{2}$ and $i<j$. The optimization evolves over 50 iterations of our routine. The initial Hamiltonian is set to $H(t)=0$ at all times. Special care must be taken since our optimal control protocol simultaneously affects the instantaneous quantum speed, $\dot{\varphi}(t)$, and its upper bound, $\Delta H(t)$. The latter varies over time for the considered time-dependent Hamiltonians. This is why we use the ratio of these quantities, namely the QSL efficiency ($\hbar\dot{\varphi}/\Delta H$), as a figure of merit.
Figure 1: QSL efficiency after geometric optimization with full Hamiltonian control: (a) Driving without QSL optimization ($p_{1}=p_{2}=0$). (b): Optimization of the driving with only the plane condition ($p_{1}=0.5,p_{2}=0$). (c): Optimization of the driving with only the phase condition ($p_{1}=0$,$p_{2}=0.1$) (d): Optimization of the driving with both conditions ($p_{1}=0.5$,$p_{2}=0.1$). In each panel, the solid line represents the QSL efficiency, i.e. ratio of the instantaneous velocity to the QSL $\hbar\dot{\phi}(t)/\Delta H(t)$, the dashed line $f_{\rm Plane}(t)$ and the dotted line $10\times f_{\rm Phase}(t)$ as a function of the rescaled time $t/t_{f}$. The 16-dimensional control vector $\mathbf{v}(t)=(\Delta_{ii}(t),\Omega_{ij}^{R}(t),\Omega_{ij}^{I}(t))$ for $(i,j)\in\{1,2,3,4\}^{2}$,$i<j$ is initialized using $\mathbf{v}^{(0)}(t)=v_{0}(1,..,1)^{T}$ with $v_{0}=0.01/t_{f}$ for $0\leq t\leq t_{f}$. Other numerical parameters: $50$ iterations, $p_{3}=10^{-4}$, $\epsilon=0.5/t_{f}$, discretization in $N_{t}=200$ times bins.
In the following, we discuss the different trade-offs at work depending on the different weights $p_{1}$, $p_{2}$. Figure 1(a) illustrates an example of optimization focusing solely on the fidelity to the target gate, with weights set as $(p_{1}=p_{2}=0)$. In this case, the routine achieves a near-perfect fidelity of $\mathcal{F}\simeq 1-10^{-4}$; however, the ratio of the quantum speed to the QSL (solid line) decreases to $70\%$ near the final time. A degradation of the QSL efficiency ($\dot{\varphi}/\Delta H$) occurs when either $f_{\rm Plane}(t)$ (dashed line) or $f_{\rm Phase}(t)$ (dotted line shows $10\times f_{\rm Phase}(t)$) begins to increase.
Figure 1(b) depicts an optimization that considers both the fidelity and the plane condition, where $f_{\rm Plane}(t)=||U_{\perp\perp}(t)||^{2}$, with weights $(p_{1}=0.5,p_{2}=0)$. The QSL efficiency ($\dot{\phi}/\Delta H(t)$) (solid line) remains above $90\%$ throughout, while $f_{\rm Plane}(t)$ (dashed line) is significantly smaller compared to the previous example. However, the final fidelity is slightly reduced, reaching $\mathcal{F}\simeq 0.95$.
Figure 1(c) corresponds to an optimization that minimizes the phase function only, with weights $(p_{1}=0,p_{2}=0.1)$. Accordingly, $f_{\rm Phase}(t)$ is smaller than in Figures 1(a,b). The QSL efficiency remains around $90\%$, though it is slightly less efficient compared to when only the plane condition is enforced (see Fig. 1(b)). The final fidelity is in this case $\mathcal{F}\simeq 0.92$.
Finally, Figure 1(d) shows an optimization that simultaneously incorporates both the plane and phase functions, with weights $(p_{1}=0.5,p_{2}=0.1)$. This approach yields the best results in terms of QSL efficiency, which stays above $95\%$ at all times. Interestingly, the plane and phase functions are smaller than when optimized separately. However, the final gate fidelity is slightly lower, with $\mathcal{F}\simeq 0.90$.
One may also consider a setup with more limited access to the Hamiltonian couplings, for example, assuming that the imaginary couplings ${\Omega_{ij}^{I}(t)}$ are unavailable. This corresponds to a scenario where Rabi pulses $\Omega_{ij}(t)$ can be applied between all sub-levels, but with the phase fixed to $\phi_{ij}(t)=0$. In this case, one can control the detunings ${\Delta_{ii}(t),1\leq i\leq 4}$, and similar results are obtained. However, if only one detuning ${\Delta_{11}(t)}$ is accessible, the quantum driving becomes less efficient, as shown in Figure 2(a-d). Even though the target quantum gate can still be achieved with very high fidelity, $\mathcal{F}\simeq 0.999$ with $p_{1}=p_{2}=0$ (Figure 2(a)), the QSL efficiency is degraded with respect to the full-control scenario depicted in Figure 1(a).
In this case, minimizing the plane condition $f_{\rm Plane}(t)$ (Figure 2(b)) results in a QSL efficiency dropping to 85%, similar to the case with full Hamiltonian control. $f_{\rm Phase}(t)$ (Figure 2(c)) under such limited access to the Hamiltonian proves counterproductive, as it reduces the QSL efficiency to 60%, compared to 70% in the absence of QSL optimization. The highest QSL efficiency is again achieved by simultaneously optimizing both the plane and phase functions, $f_{\rm Plane}(t)$ and $f_{\rm Phase}(t)$ (Figure 2(d)). In this case, the QSL efficiency reaches 90%, slightly below the QSL threshold of 95% achievable with full parameter tuning. Similarly, the final gate fidelity $\mathcal{F}\simeq 0.87$ is also slightly lower than the case of full parameter tuning.
These examples suggest that the“plane condition” associated with $f_{\rm Plane}(t)$ plays a more critical role than the“phase condition” associated with $f_{\rm Phase}(t)$. When considering these functions, an optimization involving only $f_{\rm Plane}(t)$ yields better results than one involving only $f_{\rm Phase}(t)$. Nevertheless, as expected, the optimization procedure results in better trade-offs between QSL and quantum fidelity when both functions are minimized simultaneously. The performance of the procedure improves with the number of available control parameters, as the geometric optimization routine can then explore a broader range of quantum paths.
Figure 2: QSL efficiency after geometric optimization with a limited Hamiltonian control: (a) Driving without QSL optimization ($p_{1}=p_{2}=0$). (b): Optimization of the driving with only the plane condition ($p_{1}=0.5,p_{2}=0$). (c): Optimization of the driving with only the phase condition ($p_{1}=0$,$p_{2}=0.1$) (d): Optimization of the driving with both conditions ($p_{1}=0.5$,$p_{2}=0.1$). The 7-dimensional control vector, $\mathbf{v}(t)=(\Delta_{11}(t),\Omega_{ij}^{R}(t))$ with an alphabetic ordering for $(i,j)\in\{1,2,3,4\}^{2}$ and $i<j$ is initialized using $\mathbf{v}^{(0)}(t)=v_{0}(1,1,1,1,1,1,1)^{T}$ with $v_{0}=0.01/t_{f}$ for $0\leq t\leq t_{f}$. Remaining control parameters are such that $\Delta_{ii}(t)=0$ for $i\geq 2$ and $\Omega_{ij}^{I}(t)=0$ at all times. Other numerical parameters and line conventions are identical to those of Fig. 1.
IV CONCLUSION
To conclude, we have adapted the derivation of the quantum speed limit developed by Vaidmann [3] for the evolution of single states in the context of quantum unitary operators. This derivation highlights that the quantum operators must satisfy specific geometric conditions. In particular, the time derivative of the evolution operator must always lie in a plane spanned by the initial operator and the current operator in order to saturate the MT-QSL. Since the evolution operator must also remain unitary at all times, this imposes a stringent constraint, and the MT-QSL cannot always be saturated. We have developed an optimal control algorithm to design drivings that take into account geometric constraints to approach at best the QSL. More generally, the distance of the evolution operator (or the instantaneous quantum state) from this plane provides a quantitative measure of the deviation from QSL optimality. Our method could be effectively applied to improve the QSL efficiency in a wide variety of shortcut-to-adiabaticity techniques [40]. Since the existence of a speed limit is fundamentally connected to energy and entropy exchange rates through a time-information uncertainty relation [41], this methodology extends beyond quantum systems to include, for instance, classical stochastic systems governed by Liouvillian dynamics [17, 18].
Acknowledgements We thank the support from the ANR project QuCoBEC (ANR-22-CE47-0008-02) and from the CAPES-COFECUB (20232475706P) program. F.I. acknowledges support from the Brazilian agencies CNPq (305638/2023-8), FAPERJ (210.570/2024), and INCT-IQ (465469/2014-0).
APPENDIX: DERIVATION OF THE EQUATION FOR THE ADJOINT OPERATOR DYNAMICS
Let us now put the dynamical equation for the adjoint $\mathbf{p}(t)$ in a more explicit form. From Eqs.(31,32), one finds:
$\dot{\mathbf{p}}=\frac{1}{t_{f}}\frac{\partial f_{0}(\mathbf{x}(t),\mathbf{v}(t),t)}{\partial\mathbf{x}}-\frac{\partial}{\partial\mathbf{x}}\left[\mathbf{p}(t)\cdot f(\mathbf{x}(t),\mathbf{v}(t),t)\right].$ (39)
If one folds the $2N^{2}$ column vector $\mathbf{p}(t)$ into a $2N\times N$ matrix $\mathbf{P}(t)$ (maintaining the same correspondence between $\mathbf{P}(t)$ and $\mathbf{p}(t)$ as between $\mathbf{X}(t)$ and $\mathbf{x}(t)$), one has:
$\mathbf{p}(t)\cdot f(\mathbf{x}(t),\mathbf{v}(t),t)=$
$\>{}^{T}\mathbf{P}(t)\left(\begin{array}[]{cc}H_{I}(\mathbf{v}(t))&H_{R}(\mathbf{v}(t))\\ -H_{R}(\mathbf{v}(t))&H_{I}(\mathbf{v}(t))\end{array}\right)\mathbf{X}(t).$ (42)
Equation (39) can then be recast as a differential equation for the matrix $\mathbf{P}(t)$:
$\dot{\mathbf{P}}=\frac{1}{t_{f}}\left(\begin{array}[]{c}\frac{\partial f_{0}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{R}}\\ \frac{\partial f_{0}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{I}}\end{array}\right)$ (45)
$+\frac{1}{\hbar}\left(\begin{array}[]{cc}H_{I}(\mathbf{v}(t))&H_{R}(\mathbf{v}(t))\\ -H_{R}(\mathbf{v}(t))&H_{I}(\mathbf{v}(t))\end{array}\right)\mathbf{P}(t).$ (48)
We have used the Hermiticity of the Hamiltonian, i.e. $H_{R}(\mathbf{v}(t))^{T}=H_{R}(\mathbf{v}(t))$ and $H_{I}(\mathbf{v}(t))^{T}=-H_{I}(\mathbf{v}(t))$. The second term in the dynamical equation for the adjoint vector $\mathbf{P}(t)$ is identical to the term driving the operator $\mathbf{X}(t)$, and thus corresponds to the Hamiltonian evolution for the adjoint evolution operator $U^{P}(t)=U_{R}^{P}(t)+iU_{I}^{P}(t)$ constructed from the adjoint vector (setting $\mathbf{P}(t)=(U_{R}^{P}(t),U_{I}^{P}(t))^{T}$). Summing up, one arrives at Eq. (35), with the extra driving terms associated to the derivatives of the cost function listed below (the cost associated to the Hamiltonian variance does not contribute to the dynamics of the adjoint operator):
$\frac{\partial f_{\rm Plane}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{R}}$ $=$ $-\frac{2}{d^{2}\Delta H^{2}}\left[\left({\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{R}]\right)\left(H_{R}-\frac{1}{d}{\rm Tr}[H_{R}]\mathbf{1}_{d}\right)\right.$
$\left.-\left({\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{I}]\right)H_{I}\right]-\frac{2}{d^{2}}{\rm Tr}[U_{R}]\mathbf{1}_{d},$
$\frac{\partial f_{\rm Plane}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{I}}$ $=$ $-\frac{2}{d^{2}\Delta H^{2}}\left[\left({\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{R}]\right)H_{I}\right.$
$+$ $\left.\left({\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]-\frac{1}{d}{\rm Tr}[H_{R}]{\rm Tr}[U_{I}]\right)\left(H_{R}-\frac{1}{d}{\rm Tr}[H_{R}]\mathbf{1}_{d}\right)\right]-\frac{2}{d^{2}}{\rm Tr}[U_{I}]\mathbf{1}_{d},$
$\frac{\partial f_{\rm Phase}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{R}}$ $=$ $\frac{2}{d^{2}}\sqrt{f_{\rm Phase}(\mathbf{U}(t),\mathbf{v}(t))}\left(-{\rm Tr}[U_{I}]H_{I}+{\rm Tr}[U_{R}]H_{R}+{\rm Tr}[H_{R}U_{R}-H_{I}U_{I}]\mathbf{1}_{d}-\frac{2}{d}{\rm Tr}[U_{R}]{\rm Tr}[H_{R}]\right),$
$\frac{\partial f_{\rm Phase}(\mathbf{U}(t),\mathbf{v}(t))}{\partial U_{I}}$ $=$ $\frac{2}{d^{2}}\sqrt{f_{\rm Phase}(\mathbf{U}(t),\mathbf{v}(t))}\left({\rm Tr}[U_{I}]H_{R}+{\rm Tr}[U_{R}]H_{I}+{\rm Tr}[H_{R}U_{I}+H_{I}U_{R}]\mathbf{1}_{d}-\frac{2}{d}{\rm Tr}[U_{I}]{\rm Tr}[H_{R}]\right).$ (49)